OpenStudy (anonymous):

Write an equation of the line that is perpendicular to the line 2x + 5y = -15 and contains the point (-8,3). Another one of these @Littlebird I'm going to try do it and see if I get it ;P

OpenStudy (anonymous):

First I convert the standard form of 2x + 5x = -15 into y intercept form. I get: y = -2/5x - 3 Then I input the point (-8,3) 3 = -2/5(-8) + b Simplified 3 = 3 1/5 + b Subtracted 3 1/5 from both sides Got: -1/5 = b @Littlebird @DanJS Is this correct?

OpenStudy (anonymous):

@wio @nincompoop @Jhannybean @Jesstho.-.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

You kept the slope the same

OpenStudy (jhannybean):

looks correct.

OpenStudy (anonymous):

So you are using parallel line

OpenStudy (anonymous):

OH YES, I knew I forgot something

OpenStudy (anonymous):

Parallel slope to -2/5 is 2 1/5 right?

OpenStudy (anonymous):

negative inverse

OpenStudy (anonymous):

\[2x + 5y\] will be perpendicular to \[ -5x+2y \]

OpenStudy (anonymous):

Oh, it asks for the perpendicular of the line, not the slope...

OpenStudy (anonymous):

So I have to negative inverse everything o-o

OpenStudy (anonymous):

Basically, when you have: \[ ax+by=c \]Then the perpendicular line will be a form like: \[ -bx+ay \]Does that make sense?

OpenStudy (anonymous):

In our case, we have: \[ -5x+2y = \ldots \]To find out what \(\ldots\) is, we have to plug in the point they gave us.

OpenStudy (anonymous):

(-8,3)

OpenStudy (anonymous):

So we get: \[ -5(-8)+2(3)=\ldots \]

OpenStudy (anonymous):

How did you get -5x + 2y from 2x + 5y = -15 Did you switch the variables around?

OpenStudy (anonymous):

I thought you were supposed to simplify the line from 2x + 5y = -15 into slope intercept form

OpenStudy (anonymous):

Well, the easiest way to explain it is with vectors, but you don't know vectors yet.

OpenStudy (anonymous):

I mean, you could even use \(5x-2y\) and it would work as well

OpenStudy (anonymous):

Since you're just multiplying by \(-1\).

OpenStudy (anonymous):

Another way to explain it, other than vectors, is through rotations.

OpenStudy (anonymous):

do you remember how to rotate a point?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\((x,y)\) rotated by \(90^\circ\) is going to be \((-y,x)\).

OpenStudy (anonymous):

Basically \[ \color{red}{(x,y)}\to \color{blue}{(-y,x)} \]For our line: \[ 2\color{red}x + 5\color{red}y = -15 \to 2(\color{blue}{-y})+5(\color{blue}{x})=-15 \]This simplifies to:\[ 5x-2y=-15 \]However, this line doesn't necessarily go through our point, which is why I dropped the \(-15\).

OpenStudy (anonymous):



OpenStudy (anonymous):

Anyway, since: \[ -5(-8)+2(3)=-40+6=-34 \]Our line is given by: \[ -5x+2y=-34 \]Alternatively:\[ 5x-2y=34 \]Anyway, if you want, you can put it into slope intercept form.

OpenStudy (anonymous):

@SarahLove<3 Does it make sense?

OpenStudy (anonymous):

Yeah, the way it was taught for us was to use the point slope form, just remembered that.

OpenStudy (anonymous):

Your way makes sense, perpendicular is basically a line that forms a right angle with the original line(90 degree angle) and that's what you showed in your picture with rotating a point.

OpenStudy (anonymous):

It's a bit more complex though, from what I'm learning atm ;P haha

OpenStudy (anonymous):

I got y = 5/2x + 23

OpenStudy (anonymous):

I simplified yours and got y = -5/2x + 17

OpenStudy (anonymous):

Point Slope Form Way: y - 3 = 5/2(x - -8) negative rules y - 3 = 5/2(x + 8) Property of Distribution y - 3 = 5/2x + 20 Added 3 to both sides y = 5/2x + 23

OpenStudy (anonymous):

This was from me converting 2x + 5y = -15 into y = -2/5x - 3 Then converting the slope from -2/5x into 5/2x for the negative reciprocal. Then I did the Point Slope Form Way

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

I did the math wrong. \(-5(-8) +2(3) = 40+6=46\)

OpenStudy (anonymous):

So I should have gotten:\[ -5x+2y=46\implies y=\frac 52x+23 \]

OpenStudy (anonymous):

You were right, just checked my work

OpenStudy (anonymous):

I missed something in my equation and I reworked it and I got the same thing as you

OpenStudy (anonymous):

My way is a bit different, but it isn't complex.

OpenStudy (anonymous):

Basically the trick is \(ax+by\to bx-ay\).

OpenStudy (anonymous):

However, some people just like to use formulas and stuff.

OpenStudy (anonymous):

\[ y=mx+b\implies y-mx=b\to x+my=b\implies y=-\frac 1m x+b \]

OpenStudy (anonymous):

The whole "negative inverse" \(-1/m\), is actually a result of the other method I gave.

OpenStudy (anonymous):

Wio so you mean I was right the whole time!?!?!? ;O

OpenStudy (anonymous):

No, if you keep the slope the same it is a parallel line.

OpenStudy (anonymous):

I got y = 5/2x + 23, WHEW